OpenStudy (anonymous):

Can someone help me: suppose 5 people can peel 8 baskets of apples in 3 hours. write an expression -5 people can peel 16 baskets in ? hours

OpenStudy (ali2x2):

is this multiple choice?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its fill in the blank

OpenStudy (ali2x2):

lol

OpenStudy (ali2x2):

ok so look, if 5 people can peel 8 baskets in \[3\]hours

OpenStudy (ali2x2):

then the next sentence says 5 people can peel 16 is black hours

OpenStudy (ali2x2):

basically this is 2x3

OpenStudy (ali2x2):

:)

OpenStudy (anonymous):

I'm not sure how to write an expression....

OpenStudy (anonymous):

I got 5=8x/3, so x = 15/8 for the first part. I'm not sure if the answer to the second part would be 5=30/x